OpenStudy (anonymous):

A 2 liter bottle of Pepsit costs $0.99. A gallon of milk costs $1.89. A 12 ounce can of "Food Lion" cola costs $0.20. Which fluid is the cheapest per unit?

OpenStudy (anonymous):

get all your drinks to the same unit and then calculate cost per unit. In this case I'd uses ounces as 2 of the 3 are already there. Convert 2 liters to oz using 1l = 33.814 oz. 2 l = 2 * 33.814 = 67.628 oz. We want Cost / Unit 0.99 / 67.628 = 0.01464 1.89 / 128 = 0.01477 0.20 / 12 = 0.01667 so your answer is Pepsi in the 2 liter bottle, though that's a great price for milk these days.